OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 7 times x minus 21 all over 18 minus 6 times x.?

zepdrix (zepdrix):

\[\Large \frac{7x-21}{18-6x}\]Factoring a 7 out of each term in the top,\[\Large \frac{7(x-3)}{18-6x}\]Factoring a 6 out of each term in the denominator,\[\Large \frac{7(x-3)}{6(3-x)}\]This next step is a little tricky, lemme know if it confuses you. We want to factor out a -1 from each term in the denominator,\[\Large \frac{7(x-3)}{-6(x-3)}\] See our cancellation from here? Hah, you're a duck! :D

OpenStudy (anonymous):

so > 7 from the numerator and 6 from the denominator ?

OpenStudy (anonymous):

and yes yes i am :D

zepdrix (zepdrix):

We had to do some factoring to find our `matching` factor in the numerator and denominator. The `matching` factors are the ones that will cancel out.

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

OHHHHHH thanks mate haha XD

zepdrix (zepdrix):

np c: